Casio Corp. manufactured a gold-plated, diamond-studded calculator for Justin Beiber's new job as an auditor at PW-KPM-E&D. The calculator cost $3.5 million to manufacture and Casio wants to earn a gross profit margin from the sale of the calculator of 20%. Casio leased the calculator using a 16-year lease (Justin is going for partner so he will be there a long time) with semi-annual lease payments. The first lease payment is due at the lease inception date. Casio has a required rate of return (for interest) of 6%. They expect that the calculator will be worth $.5 million at the end of the lease, but the residual value is not guaranteed. What is the lease payment amount that Casio uses for this lease?
Answer #1
Cost of calculator = $                        35,00,000
Gross profit margin = 20% on sales
Cost + Gross profit = Sale price
35,00,000+0.20X = X
X = 35,00,000/0.80
Sale price = $                        43,75,000
Less : PV of residual value $                      -77,478.70
PV of Lease payments $                  42,97,521.30
PVAF for 32 payments @ 6% =                                   14.93 "=PV(6%,32,-1,,1)" TYPE IN EXCEL
Lease payment = Sale price / PVAF for 32 payments
Lease payment = 43,75,000 / 14.93
Lease payment = $                    2,87,862.32

Please Like the solution if satisfied with the answer and if any query please mention it in comments...thanks